OpenStudy (anonymous):

A line passes through (-4, 1), perpendicular to a line whose slope is -3/2 how am I supposed to graph it anyone could help

OpenStudy (austinl):

If two lines are perpendicular, their slopes are inverse opposites. \(m_1=\dfrac{a}{b},~m_2=-\dfrac{b}{a}\) And you also have the equation of a line (slope intercept form), \(y=mx+b\)

OpenStudy (anonymous):

Ok

OpenStudy (austinl):

Did that make sense?

OpenStudy (anonymous):

Yea kinda

OpenStudy (anonymous):

So if I plug it in like this... 1=-3/2x +(-4)

OpenStudy (austinl):

Wrong slope, \(m_1=-\dfrac{3}{2},~m_2=~?\)

OpenStudy (austinl):

And you plugged in the x value into the "b" spot, you plug it in for "x".

OpenStudy (anonymous):

How do u get to kno the slope of the other line?

OpenStudy (austinl):

I told you how to find it up above. Inverse opposite of the given slope.

OpenStudy (anonymous):

Oooh so if the slope is negative I change it to inverse its positive?

OpenStudy (anonymous):

No its not making any sense now explain in a different way if u can

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

@hubbabubba : When tow line are perpendicular we have : m.m2=-1 So the slope is 2/3 And we have : y=2/3x But what is the (-4,1) ? Look : -4/1=-4 So we should be have d too . Now can u answer d ?:)

OpenStudy (anonymous):

Where did u get d from???

OpenStudy (anonymous):

d is the piont the line and y .

OpenStudy (anonymous):

Wait ... What is the (-4,1)in the question ???

OpenStudy (anonymous):

Its a line that passes thro tha point

OpenStudy (anonymous):

It s a point ?

OpenStudy (anonymous):

@hubbabubba

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So its ok ! Now u should use this formula : y-y2=m(x-x2) X1=-4 Y2=1 OK?:)

OpenStudy (anonymous):

K

OpenStudy (anonymous):

So its gonna be

OpenStudy (anonymous):

1 - y2 = -3/2 (-4 - x2) right?

OpenStudy (anonymous):

Excelent but I said to you : x2=-4 y2=1 2 is just a number no *2 ...

OpenStudy (anonymous):

Ik it x2

OpenStudy (anonymous):

Y x2 n y2 Y not x1 n y1

OpenStudy (anonymous):

No diffrence 1or2 are just name . you should get the things true in formula means get the (-4,1)of point than y and x not affter them ! Because your answer will be fall :) Now answer again :)

OpenStudy (anonymous):

Y - 1 = -3/2 ( x + 4 ) right?

OpenStudy (anonymous):

Exelent !:)

OpenStudy (anonymous):

Then y = -3/2x + 6 right? As in y=mx +b

OpenStudy (anonymous):

But the slope was 2/3 forgot ?

OpenStudy (anonymous):

Look : y-1=2/3(x+4) OK ??!:)

OpenStudy (austinl):

A line passes through (-4, 1), perpendicular to a line whose slope is -3/2. \(m_1=-\dfrac{a}{b}=-\dfrac{3}{2}\) \(m_2=-(-\dfrac{b}{a})=-(-\dfrac{2}{3})=\dfrac{2}{3}\) Then we have the equation of a line, \(y=mx+b\) With the point, \((-4,1)=(x,y)\) which we then plug into the equation to solve for the y intercept b. \(y=mx+b\rightarrow1=(-4)(\dfrac{2}{3})+b\) \(b=\dfrac{11}{3}\) So therefore we have, \(y=\dfrac{2}{3}x+\dfrac{11}{3}\)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

@hubbabubba :Understand and got it ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

did u get my formula ?

OpenStudy (anonymous):

How did u get 11/3 ? @ austinL

OpenStudy (anonymous):

And no ur making it complicated for me btw @E.ali

OpenStudy (anonymous):

How did u get 11/3 ? @ austinL

OpenStudy (anonymous):

@austinL

OpenStudy (austinl):

You simply solve for b in that equation.

OpenStudy (anonymous):

Dont think about 11/3 . I m not sure about it :) Can you write it by the formula ?

OpenStudy (austinl):

\(y=mx+b→1=(−4)(\dfrac{2}{3})+b\)

OpenStudy (anonymous):

Yea i did that

OpenStudy (anonymous):

Wat formula are u talking abt? Im taking algebra 2 here

OpenStudy (anonymous):

@austinL : I think I m answering friend !:)

OpenStudy (anonymous):

@hubbabubba :about about it :y-1=2/3(x+4) OK?

OpenStudy (anonymous):

For the 3rd time I got the equation I solved it but it didnt give me 11/3

OpenStudy (anonymous):

So u multiply ( -4 ) × 2/3 right?

OpenStudy (austinl):

\(-4\times(\dfrac{2}{3})=-\dfrac{8}{3}\) \(1+\dfrac{8}{3}=\dfrac{3}{3}+\dfrac{8}{3}=\dfrac{8+3}{3}=\dfrac{11}{3}\)

OpenStudy (anonymous):

Oh k makes sense now thnx austin

OpenStudy (austinl):

No problemo.

OpenStudy (anonymous):

Ok then so how do I graph the equation n the point?